# Mean-Field Microcanonical Gradient Descent

### v1

@@ 5.2 Financial data
We evaluate the model on real financial data, namely the S&P 500 index [^1] Yahoo Finance https://finance.yahoo.com/quote/%5EGSPC/history, and five and ten year synthetic EUR and USD government bonds [^2] Sveriges riksbank (Swedish Central Bank) https://www.riksbank.se/en-gb/statistics/ quoted in yield. For the equity index, daily log-returns are generated, while regular daily returns are used for the rates. To compute the target energy, $2^{12}$ points were used for S&P 500 and USD rates, and $2^{11}$ for EUR rates, corresponding to approximately 16 and 8 years worth of daily data, respectively. For generation, samples of length $2^{10}$, or about 4 years, were produced.
We use stylized features in the energy function, consisting of variance, autocovariance at lag 1, and autocovariance of the squared process for lags 1–20 — statistics that are of interest for financial time series. In Appendix C, we provide results for the scattering covariance as energy. We qualitatively compare the models based on these statistics, as well as marginal distributions, with results displayed in Figure 6 (S&P 500) and Appendix C (rates data). As reference, we also include a GARCH(1,1)11(1,1)( 1 , 1 ) model with AR$(1)$ mean process and Student’s t innovations, using maximum likelihood parameters fitted using the Python ARCH package (Sheppard, 2024). The same general behavior is observed here as for the AR$(1)$ example, namely that the MF–MGDM counteracts the overfitting while still producing a good fit of the statistics that the model is conditioned on. The marginal fit, however, which except for the variance is not explicitly optimized for, becomes slightly worse, although it still manages to capture some of the heavy tails.
@@ 6 Conclusions
The MGDM provides efficient sampling of high-dimensional distributions, but can suffer from a significant loss of entropy. Propagating too far into the descent is shown to overfit to the target energy that the model is conditioned on, meaning that the variance of the energy for the model is much too small as for what to expect from true distributions. Regularizing by early stopping in the descent mitigates this issue somewhat, but at the price of a worse fit to the true distribution and a larger bias from the initial distribution. The mean-field regularization of the model leverages parallel sampling to mitigate the problem, improving the rate at which entropy is lost without a significant impact on the likelihood fit. Future work will explore better initial distributions and more sophisticated update steps. These will in turn open the door to considering forward KL divergence metrics, removing the need for access to the likelihood of the target distribution.
Impact Statement
This paper presents work whose general goal is to advance the field of machine learning. As such, there are many potential societal consequences of our work, none which must be specifically highlighted here.
Acknowledgements
@@
Appendix A Appendix – Computing the Jacobian determinant in MF-MGDM
Denote $\tilde{g}^{(n)}$ as the update corresponding to particle $x^{(n)}$:
$\tilde{g}^{(n)}(\bm{x})=x^{(n)}-\gamma\sum_{k=1}^{K}\nabla\Phi_{k}(x^{(n)})\bar{\Phi}_{k}(\bm{x}).$
@@ Appendix A Appendix – Computing the Jacobian determinant in MF-MGDM
$\partial_{x^{(m)}_{j}}\tilde{g}^{(n)}_{i}(\bm{x})$ $=\delta_{m,n}\delta_{i,j}-\gamma\sum_{k}\partial_{x^{(m)}_{j}}\left(\partial_{x^{(n)}_{i}}\Phi_{k}(x^{(n)})\cdot\bar{\Phi}_{k}(\bm{x})\right)$
$=\delta_{m,n}\delta_{i,j}-\gamma\sum_{k}\left(\delta_{m,n}\partial_{x^{(n)}_{j}}\partial_{x^{(n)}_{i}}\Phi_{k}(x^{(n)})\cdot\bar{\Phi}_{k}(\bm{x})+\frac{1}{N}\partial_{x^{(n)}_{i}}\Phi_{k}(x^{(n)})\cdot\partial_{x^{(m)}_{j}}\Phi_{k}(x^{(m)})\right),$
@@
$J_{\tilde{g}^{(n)}}(x^{(m)})=\delta_{m,n}\cdot\left(I-\gamma\sum_{k}H_{\Phi_{k}}(x^{(n)})\bar{\Phi}_{k}(\bm{x})\right)-\frac{\gamma}{N}J_{\Phi}^{\top}(x^{(n)})J_{\Phi}(x^{(m)}).$
Define the block matrices
$\mathcal{J}_{\Phi}(\bm{x})=\begin{bmatrix}J_{\Phi}(x^{(1)})&\cdots&J_{\Phi}(x^{(N)})\end{bmatrix}$
and
$\mathcal{H}_{\Phi_{k}}(\bm{x})=\operatorname{diag}\left\{H_{\Phi_{k}}(x^{(n)})\right\}_{n=1}^{N}=\begin{bmatrix}H_{\Phi_{k}}(x^{(1)})&&\\ &\ddots&\\ &&H_{\Phi_{k}}(x^{(N)})\end{bmatrix}.$
Then, the entire Jacobian of $\tilde{g}$ can be expressed as
$J_{\tilde{g}}(\bm{x})=I-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}(\bm{x})\bar{\Phi}_{k}(\bm{x})-\frac{\gamma}{N}\mathcal{J}_{\Phi}^{\top}\mathcal{J}_{\Phi}(\bm{x}).$
@@
$I-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}\bar{\Phi}_{k}$
@@
$\det J_{\tilde{g}}$ $=\det\left(I-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}\bar{\Phi}_{k}-\frac{\gamma}{N}\mathcal{J}_{\Phi}^{\top}\mathcal{J}_{\Phi}\right)$
$=\det\left(I-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}\bar{\Phi}_{k}\right)\det\left(I-\frac{\gamma}{N}\mathcal{J}_{\Phi}\left(I-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}\bar{\Phi}_{k}\right)^{-1}\mathcal{J}_{\Phi}^{\top}\right)$
$=\det\operatorname{diag}\left\{\left(I-\gamma\sum_{k}H_{\Phi_{k}}^{(n)}\bar{\Phi}_{k}\right)\right\}_{n}\det\left(I-\frac{\gamma}{N}\mathcal{J}_{\Phi}\operatorname{diag}\left\{\left(I-\gamma\sum_{k}H_{\Phi_{k}}^{(n)}\bar{\Phi}_{k}\right)^{-1}\right\}_{n}\mathcal{J}_{\Phi}^{\top}\right)$
$=\prod_{n}\det\left(I-\gamma\sum_{k}H_{\Phi_{k}}^{(n)}\bar{\Phi}_{k}\right)\det\left(I-\gamma\frac{1}{N}\sum_{n}J_{\Phi}^{(n)}\left(I-\gamma\sum_{k}H_{\Phi_{k}}^{(n)}\bar{\Phi}_{k}\right)^{-1}\left(J_{\Phi}^{(n)}\right)^{\top}\right).$
Appendix B Appendix – Projected Gradient Descent
@@
Appendix C Appendix – Additional Plots
@@
Figure 8: Autocorrelations and marginal histograms as in Figure 6, here using the scattering covariance with phase shifts as described in Section 5.1, but with PCA components now computed using samples from a GARCH process. Since the autocorrelations are not explicitly included in the energy function, the fit is worse. For MGDM, the statistics are not as concentrated as in Figures 6 and 7. The MF–MGDM still provide some improvements, most noticeable in the ACF.
@@
Figure 10: Autocorrelations and marginal histograms with scattering covariance energy as in Figure 8, the initial distribution now coming from a GARCH process as opposed to Gaussian white noise in Figure 8(a). The fit is better for both MGDM and MF-MGDM compared to Fig. 8(a). Together with Figure 9, this shows that the shortcomings in Fig. 8(a) are due to the microcanonical set being too large (i.e., additional momentum constraints are necessary), rather than issues with the decent failing to transport the initial samples to the microcanonical set.

### v2

@@ 5.2 Financial data
Figure 6: S&P 500 realization and a randomly picked generated samples, showing a half-year window.
We evaluate the model on real financial data, namely the S&P 500 index [^1] Yahoo Finance https://finance.yahoo.com/quote/%5EGSPC/history (Terms), as well as five- and ten-year synthetic EUR and USD government bonds [^2] Sveriges riksbank (Swedish Central Bank) https://www.riksbank.se/en-gb/statistics/ (Terms) quoted in yield. For the equity index, daily log-returns are generated, while regular daily returns are used for the rates. We use $2^{12}$ points ($\sim$ 16 years of daily data) for S&P 500 and USD rates, and $2^{11}$ ($\sim$ 8 years) for EUR rates, which we divide into four samples of equal length. The energy used for generation is estimated on the first sample, and the remaining three are used for validation.
Figure 7: S&P 500 autocorrelation, autocorrelation of the square of the signal, and marginal histogram. The top two rows are violin plots illustrating the empirical marginal density of the statistics for the models, with whiskers indicating min and max. In the top two rows, the statistics of the true sample used for the target energy $\alpha$ is in black and the validation samples are thin gray. In the bottom row, the same holds, except the joint histogram for the validation data is shown. The dashed gray line is a Gaussian with moments matched to the true data. Generation time was equal for MGDM and MF–MGDM.
As energy function $\Phi$ we use statistics of interest for financial time series, namely variance, autocovariance at lag 1, and autocovariance of the squared process for lags 1–20. (In Appendix D, we also provide results for the scattering covariance as energy.) Realizations of the models conditioned on S&P 500 data are displayed in Figure 6 together with a slice of the validation data. As reference, we also include a GARCH(1,1)11(1,1)( 1 , 1 ) model with AR$(1)$ mean process and Student’s t innovations, using maximum likelihood parameters fitted using the Python ARCH package [33, v6.2] (University of Illinois/NCSA Open Source License). All three models evaluated provide samples that are qualitatively similar to the original signal. In addition, we compare the statistics included in $\Phi$ and the marginal histograms, with results displayed in Figure 7 (S&P 500) and Figure 8 in Appendix D (rates data). The same general behavior is observed here as for the AR$(1)$ example, namely that the MF–MGDM counteracts the overfitting while still producing a good fit of the statistics that the model is conditioned on, comparable to GARCH. The marginal fit, however, is superior for GARCH, with MF–MGDM becoming slightly worse than MGDM. As a remedy, the energy function could be extended to include more sophisticated statistics to incorporate the heavy tails in both gradient models. Finally, we estimated the entropy for the two microcanonical approximations to –48 800 for MGDM and +1 200 for MF–MGDM, in relation to +1 450 for Gaussian white noise, indicating improved performance with the mean-field approach.
6 Limitations
First, we emphasize the stationarity assumption of the time series. Next, the MGDM requires the energy function $\Phi$ to be differentiable so it is not straightforward to include e.g. order statistics constraints. As far as we know, there is presently no modification of the MGDM which allows for a stable way of inverting the descent in order to be able to compute forward KL in the usual case where the true distribution is not known, forcing only qualitative evaluation of performance on real-world data. Note also that in this case, any (differentiable) evaluation statistic that is of interest can be included in $\Phi$, which in turn risks weakening the merit of the evaluation akin to Goodhart’s law. Finally, although the width $\varepsilon$ of $\Omega_{\varepsilon}$ is important for a good KL fit, exactly how to tune this parameter is left for future work.
@@ 7 Conclusions
The MGDM provides efficient sampling of high-dimensional distributions, but can suffer from a significant loss of entropy. Propagating too far into the descent is shown to overfit to the target energy that the model is conditioned on, meaning that the variance of the energy for the model is much too small as for what to expect from true distributions. Regularizing by early stopping in the descent mitigates this issue somewhat, but at the price of a worse fit to the true distribution and a larger bias from the initial distribution. The mean-field regularization of the model in the form of MF–MGDM leverages parallel sampling to mitigate the problem, improving the rate at which entropy is lost without a significant impact on the likelihood fit. Future work will explore better initial distributions and more sophisticated update steps. These will in turn open the door to considering forward KL divergence metrics, removing the need for access to the likelihood of the target distribution.
Acknowledgments and Disclosure of Funding
@@
Appendix A Appendix – Computing the Jacobian determinant in MF–MGDM
Without loss of generality, assume $\alpha=0$ (otherwise, we simply redefine $\Phi(x)$ to be $\Phi(x)-\alpha$). Denote $\overline{g}^{(n)}$ as the update corresponding to particle $x^{(n)}$:
$\overline{g}^{(n)}(\bm{x})=x^{(n)}-\gamma\sum_{k=1}^{K}\nabla\Phi_{k}(x^{(n)})\overline{\Phi}_{k}(\bm{x}).$
@@ Appendix A Appendix – Computing the Jacobian determinant in MF–MGDM
$\partial_{x^{(m)}_{j}}\overline{g}^{(n)}_{i}(\bm{x})$ $=\delta_{m,n}\delta_{i,j}-\gamma\sum_{k}\partial_{x^{(m)}_{j}}\left(\partial_{x^{(n)}_{i}}\Phi_{k}(x^{(n)})\cdot\overline{\Phi}_{k}(\bm{x})\right)$
$=\delta_{m,n}\delta_{i,j}-\gamma\sum_{k}\left(\delta_{m,n}\partial_{x^{(n)}_{j}}\partial_{x^{(n)}_{i}}\Phi_{k}(x^{(n)})\cdot\overline{\Phi}_{k}(\bm{x})+\frac{1}{N}\partial_{x^{(n)}_{i}}\Phi_{k}(x^{(n)})\cdot\partial_{x^{(m)}_{j}}\Phi_{k}(x^{(m)})\right),$
@@
$J_{\overline{g}^{(n)}}(x^{(m)})=\delta_{m,n}\cdot\left(\text{I}_{d}-\gamma\sum_{k}H_{\Phi_{k}}(x^{(n)})\overline{\Phi}_{k}(\bm{x})\right)-\frac{\gamma}{N}J_{\Phi}^{\top}(x^{(n)})J_{\Phi}(x^{(m)}),$
where $\text{I}_{d}$ is the $d\times d$ identity matrix. Recall the concatenation (10) of the Jacobians,
$\mathcal{J}_{\Phi}(\bm{x})=\begin{bmatrix}J_{\Phi}(x^{(1)})&\cdots&J_{\Phi}(x^{(N)})\end{bmatrix},$
and define the block-diagonal matrix
$\mathcal{H}_{\Phi_{k}}(\bm{x})=\operatorname{diag}\left\{H_{\Phi_{k}}(x^{(n)})\right\}_{n=1}^{N}=\begin{bmatrix}H_{\Phi_{k}}(x^{(1)})&&\\ &\ddots&\\ &&H_{\Phi_{k}}(x^{(N)})\end{bmatrix}.$ (12)
Then, the entire Jacobian of $\overline{g}$ can be expressed as
$J_{\overline{g}}(\bm{x})=\text{I}_{Nd}-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}(\bm{x})\overline{\Phi}_{k}(\bm{x})-\frac{\gamma}{N}\mathcal{J}_{\Phi}^{\top}(\bm{x})\mathcal{J}_{\Phi}(\bm{x}).$ (13)
@@
$\text{I}_{Nd}-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}\overline{\Phi}_{k}$
@@
$\det J_{\overline{g}}$ $=\det\left(\text{I}_{Nd}-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}\overline{\Phi}_{k}-\frac{\gamma}{N}\mathcal{J}_{\Phi}^{\top}\mathcal{J}_{\Phi}\right)$
$=\det\left(\text{I}_{Nd}-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}\overline{\Phi}_{k}\right)\det\left(\text{I}_{K}-\frac{\gamma}{N}\mathcal{J}_{\Phi}\left(\text{I}_{Nd}-\gamma\sum_{k}\mathcal{H}_{\Phi_{k}}\overline{\Phi}_{k}\right)^{-1}\mathcal{J}_{\Phi}^{\top}\right)$
$=\det\operatorname{diag}\left\{\left(\text{I}_{d}-\gamma\sum_{k}H_{\Phi_{k}}^{(n)}\overline{\Phi}_{k}\right)\right\}_{n}\det\left(\text{I}_{K}-\frac{\gamma}{N}\mathcal{J}_{\Phi}\operatorname{diag}\left\{\left(\text{I}_{d}-\gamma\sum_{k}H_{\Phi_{k}}^{(n)}\overline{\Phi}_{k}\right)^{-1}\right\}_{n}\mathcal{J}_{\Phi}^{\top}\right)$
$=\prod_{n}\det\left(\text{I}_{d}-\gamma\sum_{k}H_{\Phi_{k}}^{(n)}\overline{\Phi}_{k}\right)\det\left(\text{I}_{K}-\gamma\frac{1}{N}\sum_{n}J_{\Phi}^{(n)}\left(\text{I}_{d}-\gamma\sum_{k}H_{\Phi_{k}}^{(n)}\overline{\Phi}_{k}\right)^{-1}\left(J_{\Phi}^{(n)}\right)^{\top}\right).$
Appendix B Appendix – Proof of Theorem 4.1
As in previous Section A, we assume without loss of generality that $\alpha=0$.
From (8) we get
$H(\overline{q}_{T}^{N})=-\mathbb{E}_{\overline{q}_{T}^{N}}[\log\overline{q}_{T}^{N}(\bm{X})]$ $=-\mathbb{E}_{\overline{q}_{0}^{N}}\left[\log\overline{q}_{0}^{N}(\bm{X})-\sum_{t=0}^{T-1}\log|\det J_{\overline{g}}(\overline{g}_{t}(\bm{X}))|\right]$
$=H(\overline{q}_{0}^{N})+\sum_{t=0}^{T-1}\mathbb{E}_{\overline{q}_{t}^{N}}[\log|\det J_{\overline{g}}(\bm{X})|.]$ (14)
so we want to lower-bound $\log|\det J_{\overline{g}}|$. By (13) we see that we can write $J_{\overline{g}}(\bm{x})$ on the form $\text{I}-\gamma A$. We have
$\left.\frac{d}{d\gamma}\det(\text{I}-\gamma A)\right|_{\gamma=0}=-\det(\text{I})\operatorname{Tr}(\text{I}^{-1}A)=-\operatorname{Tr}A,$
which yields the Taylor approximation
$\det(\text{I}-\gamma A)=1-\gamma\operatorname{Tr}A+\operatorname{\mathcal{O}}(\gamma^{2}).$
This, together with the lower bound for the logarithm
$\log(1-x)\geq-2x$
for $x\in[0,\frac{3}{4}]$, results in the lower bound (suppressing the argument $(\bm{x})$)
$\log|\det J_{\overline{g}}|\geq-2\gamma\left|\operatorname{Tr}\left(\sum_{k}\mathcal{H}_{\Phi_{k}}\overline{\Phi}_{k}+\frac{1}{N}\mathcal{J}_{\Phi}^{\top}\mathcal{J}_{\Phi}\right)\right|+\operatorname{\mathcal{O}}(\gamma^{2})$ (15)
for $\gamma$ small enough. Thus, we seek an upper bound to
$\left|\operatorname{Tr}\left(\sum_{k}\mathcal{H}_{\Phi_{k}}\overline{\Phi}_{k}+\frac{1}{N}\mathcal{J}_{\Phi}^{\top}\mathcal{J}_{\Phi}\right)\right|\leq\sum_{k}|\operatorname{Tr}(\mathcal{H}_{\Phi_{k}})\overline{\Phi}_{k}|+\frac{1}{N}|\operatorname{Tr}(\mathcal{J}_{\Phi}^{\top}\mathcal{J}_{\Phi})|.$ (16)
The Lipschitz assumption on $\Phi$ yields $\|J_{\Phi}(x)\|_{2}\leq\beta$ for all $x$, so that for any particle (here suppressing the argument $(x^{(i)})$),
$\operatorname{Tr}(J_{\Phi}^{\top}J_{\Phi})=\operatorname{Tr}(J_{\Phi}J_{\Phi}^{\top})=\sum_{k}\lambda_{k}(J_{\Phi}J_{\Phi}^{\top})\leq K\lambda_{\max}(J_{\Phi}J_{\Phi}^{\top})=K\|J_{\Phi}^{\top}\|_{2}^{2}=K\|J_{\Phi}\|_{2}^{2}\leq K\beta^{2},$
whereby the second term of (16) becomes
$\frac{1}{N}\operatorname{Tr}(\mathcal{J}_{\Phi}^{\top}(\bm{x})\mathcal{J}_{\Phi}(\bm{x}))=\frac{1}{N}\sum_{i=1}^{N}\operatorname{Tr}(J_{\Phi}^{\top}(x^{(i)})J_{\Phi}(x^{(i)}))\leq\frac{1}{N}NK\beta^{2}=K\beta^{2}.$ (17)
Similarly, the Lipschitz assumption on $\nabla\Phi$ together with symmetry of $H$ implies $\|H_{\Phi_{k}}(x)\|_{2}=|\lambda|_{\max}(H_{\Phi_{k}}(x))\leq\eta$ for all $k$ and $x$, and in turn,
$\sum_{k}|\operatorname{Tr}(H_{\Phi_{k}})\overline{\Phi}_{k}|\leq\sum_{k}d|\lambda|_{\max}(H_{\Phi_{k}})|\overline{\Phi}_{k}|\leq d\eta\|\overline{\Phi}\|_{1}\leq d\eta\sqrt{K}\|\overline{\Phi}\|_{2}.$
Thus, the first term of (16) becomes
$\sum_{k}|\operatorname{Tr}(\mathcal{H}_{\Phi_{k}}(\bm{x}))\overline{\Phi}_{k}(\bm{x})|=\sum_{k}\left|\sum_{i=1}^{N}\operatorname{Tr}(H_{\Phi_{k}}(x^{(i)}))\overline{\Phi}_{k}(\bm{x})\right|\leq Nd\eta\sqrt{K}\|\overline{\Phi}(\bm{x})\|_{2}.$ (18)
Inserting (17) and (18) into (16), we see that the $\log|\det J_{\overline{g}}|$ bound (15) becomes
$\log|\det J_{\overline{g}}(\bm{x})|\geq-2\gamma\left(Nd\eta\sqrt{K}\|\overline{\Phi}(\bm{x})\|_{2}+K\beta^{2}\right)+\operatorname{\mathcal{O}}(\gamma^{2}).$
Hence, the lower bound on the entropy rate, up to second order terms in $\gamma$, becomes
$(Nd)^{-1}H(\overline{q}_{T}^{N})$ $=(Nd)^{-1}H(\overline{q}_{0}^{N})-2\gamma\left(\eta\sqrt{K}\sum_{t=0}^{T-1}\mathbb{E}_{\overline{q}_{t}^{N}}\|\overline{\Phi}(\bm{X})\|_{2}+\frac{K}{Nd}\beta^{2}T\right).$
Appendix C Appendix – Projected gradient descent
@@
Another aspect of the projected version of MGDM is the choice of initial measure. If the support of the marginal distribution is all of $\mathbb{R}$, the maximum entropy distribution conditioned on the first two moments is the Gaussian. Thus, in this case, the MGDM is initialized with Gaussian white noise. For the CIR process, the support of the marginal distribution is $\mathbb{R}^{+}$, and, given that it exists, the corresponding maximum entropy distribution is either the exponential (if the mean and standard deviation are equal) or the truncated Gaussian [34].
Appendix D Appendix – Additional plots
@@
Figure 9: Autocorrelations and marginal histograms as in Figure 7, here using the scattering covariance with phase shifts as described in Section 5.1, but with PCA components now computed using samples from a GARCH process. Since the autocorrelations are not explicitly included in the energy function, the fit is worse. For MGDM, the statistics are not as concentrated as in Figures 7 and 8. The MF–MGDM still provide some improvements, most noticeable in the ACF.
@@
Figure 11: Autocorrelations and marginal histograms with scattering covariance energy as in Figure 9, the initial distribution now coming from a GARCH process as opposed to Gaussian white noise in Figure 9(a). The fit is better for both MGDM and MF–MGDM compared to Fig. 9(a). Together with Figure 10, this shows that the shortcomings in Fig. 9(a) are due to the microcanonical set being too large (i.e., additional moment constraints are necessary), rather than issues with the descent failing to transport the initial samples to the microcanonical set.